# On the Learning with Augmented Class via Forests

### v1

@@ Abstract
Decision trees and forests have achieved successes in various real applications, most working with all testing classes known in training data. In this work, we focus on learning with augmented class via forests, where an augmented class may appear in testing data yet not in training data. We incorporate information of augmented class into trees’ splitting, i.e., a new splitting criterion, called augmented Gini impurity, is introduced to exploit some unlabeled data from testing distribution. We then develop the approach named Learning with Augmented Class via Forests (LACForest), which constructs shallow forests based on the augmented Gini impurity and then splits forests with pseudo-labeled augmented instances for better performance. We also develop deep neural forests with a novel optimization objective based on our augmented Gini impurity, so as to utilize the representation power of neural networks for forests. Theoretically, we present the convergence analysis for augmented Gini impurity, and finally conduct experiments to verify the effectiveness of our approaches. The code is available at https://github.com/nju-xuf/LACForest/.
@@ 1 Introduction
How to handle distribution change has become an important problem in a non-stationary learning environment Zhou (2022), and recent years have witnessed increasing attentions with various real applications Gama et al. (2014); Geng et al. (2021); Wang et al. (2024). This work focuses on learning with augmented class, that is, the class distribution changes and an augmented class unseen in the training data may emerge during the testing process Da et al. (2014). Here, we take the object recognition in autonomous driving for example: New objects may appear on roads yet not in the historical labeled data, and a reliable learning system should make good predictions over both known classes and augmented class.
Various approaches have been developed for learning with augmented class. Da et al. Da et al. (2014) first named this problem and adjusted decision boundaries for augmented class under the low-density assumption. Bendale and Boult Bendale and Boult (2016) and Rudd et al. Rudd et al. (2017) estimated the probability of an instance belonging to augmented class on the basis of extreme value theory. Mendes Júnior et al. Mendes Júnior et al. (2017) detected augmented class via nearest neighbor method, and Liu et al. Liu et al. (2018) presented PAC guarantees for the detection of augmented class. Some generative networks have also been applied to learn augmented class Ge et al. (2017); Neal et al. (2018); Chen et al. (2021). Zhang et al. Zhang et al. (2020) gave an unbiased risk estimation by exploiting unlabeled data, and Shu et al. Shu et al. (2023) generalized such approach to arbitrary loss functions.
Decision trees and forests have achieved great successes in various real applications with strong capabilities in handling discrete features and exploring local regions Cutler et al. (2007); Qi (2012); Grinsztajn et al. (2022); Costa and Pedreira (2023), mostly working with the assumption that training and testing data come from the same distribution. For tree models, previous studies only detected augmented class from labeled data of known classes Mu et al. (2017); Liu et al. (2018), and it still remains open to exploit unlabeled data from testing distribution to build trees for learning with augmented class.
This work tries to incorporate some useful information of augmented class into the construction of forests, and our main contributions could be summarized as follows:
  • We introduce a new splitting criterion, called augmented Gini impurity, which incorporates useful information of augmented class into trees’ splitting from unlabeled data. We then develop the approach named Learning with Augment Class via Forests (LACForest), which firstly constructs shallow forests based on augmented Gini impurity and further splits forests with pseudo-labeled augmented instances for better performance.
@@
  • Theoretically, we present the convergence analysis of our augmented Gini impurity w.r.t. both decision trees and deep neural trees. We finally conduct experiments to validate the effectiveness of our proposed methods, which perform better over the classification of all classes as well as the detection of augmented class, in contrast to existing methods on learning with augmented class.
The rest of this work is organized as follows. Section 2 presents preliminaries of this work. Section 3 introduces our augmented Gini impurity and LACForest approach. Section 4 presents our deep neural approach. Section 5 conducts extensive experiments. Section 6 concludes with future works.
@@ 2 Preliminaries
We focus on the semi-supervised setting Da et al. (2014); Liu et al. (2018); Zhang et al. (2020); Shu et al. (2023), and the goal is to learn a function $f\colon\mathcal{X}\rightarrow\mathcal{Y}$ from labeled data $S_{l}$ and unlabeled data $S_{u}$ with
$S_{l}=\{(\bm{x}^{l}_{1},y_{1}),\ldots,(\bm{x}^{l}_{n_{l}},y_{n_{l}})\}\text{ and }S_{u}=\{\bm{x}^{u}_{1},\ldots,\bm{x}^{u}_{n_{u}}\}\ ,$
where each labeled example $(\bm{x}^{l}_{i},y_{i})$ is drawn i.i.d. from $\mathcal{D}_{\text{kc}}$, and unlabeled instance $\bm{x}^{u}_{j}$ is sampled i.i.d. from $\mathcal{D}_{\mathcal{X}}$.
Let $\mathbb{I}[\cdot]$ be the indicator function, which returns $1$ if the input is true and 00 otherwise. For integer $k>0$, we write $[k]=\{1,2,\cdots,k\}$, and denote by $\Delta_{k}$ the $k$-dimensional simplex. For $z\in\mathbb{R}$, we denote by $\lfloor z\rfloor$ the largest integer no more than $z$, and define $(z)_{+}=\max(0,z)$. Let $|A|$ be the cardinality of set $A$.
@@ 3 Our LACForest approach
This section proposes the LACForest approach for learning with augmented class based on random forests, and the core idea is to introduce a new splitting criterion, augmented Gini impurity, to incorporate potential augmented instances into the process of tree splitting.
@@ 3.1 Augmented Gini impurity
Denote by $n_{\mathcal{C},l}=|S_{\mathcal{C},l}|$ and $n_{\mathcal{C},u}=|S_{\mathcal{C},u}|$. We define the augmented Gini impurity as a splitting criterion as follows:
@@
$\vartheta_{\mathcal{C},k}(S_{l},S_{u})=(1-\vartheta_{\mathcal{C},\kappa+1}(S_{l},S_{u}))\sum\limits_{(\bm{x},y)\in S_{l}}\frac{\mathbb{I}[\bm{x}\in\mathcal{C},y=k]}{\max(1,n_{\mathcal{C},l})}\ .$
Figure 1: An intuitive illustration for the difference between our criterion and original Gini impurity over a 2-dimensional synthetic dataset, due to the consideration of augmented instances.
This definition is essentially similar to the original Gini impurity Breiman (1984) except for an additional term $\vartheta_{\mathcal{C},\kappa+1}(S_{l},S_{u})$, which aims to incorporate augmented class into impurity measure. Figure 1 shows an example on a synthetic dataset, and our augmented Gini impurity could take potential augmented data into consideration compared with Gini impurity. This is because $\vartheta_{\mathcal{C},\kappa+1}(S_{l},S_{u})$ estimates the proportion of augmented instances in $\mathcal{C}$ w.r.t. distribution $\mathcal{D}$ based on Assumption 1 with following guarantee:
@@
For instance space $\mathcal{C}\subseteq\mathcal{X}$ and $\delta\in(0,1)$, the following holds with probability at least $1-\delta$ over $S_{l}$ and $S_{u}$
@@
where $\tilde{\bm{y}}\in\mathbb{R}^{\kappa+1}$ is the one-hot encoding of $y\in[\kappa+1]$. Intuitively, $\mathcal{L}^{*}_{\mathcal{C}}$ shows the minimum expected squared loss by giving the same prediction for all instances in space $\mathcal{C}$. Based on Lemma 1 and Assumption 1, we have
@@
For instance space $\mathcal{C}\subseteq\mathcal{X}$ and $\delta\in(0,1)$, the following holds with probability at least $1-\delta$ over $S_{l}$ and $S_{u}$
@@
This theorem shows that our augmented Gini impurity can be viewed as a good estimation of $\mathcal{L}^{*}_{\mathcal{C}}$, and we could get the convergence rate $O(1/\sqrt{\min\{n_{l},n_{u}\}})$. The detailed proof is given in Appendix A.2, which involves the equivalence between splitting criterions and loss functions, as well as the convergence of $\vartheta_{\mathcal{C},\kappa+1}(S_{l},S_{u})$ in Lemma 1.
@@ 3.2 Construction of LACForest
A tree model can be constructed by partitioning instance space into disjoint rectangular nodes recursively. Given node $\mathcal{C}\subseteq\mathcal{X}$, splitting feature $j\in[d]$ and splitting point $a\in\mathbb{R}$, we denote the left and right children, respectively, as
$\mathcal{C}^{l}_{j,a}=\{\bm{x}\in\mathcal{C}\colon x_{j}\leq a\}\ \text{ and }\ \mathcal{C}^{r}_{j,a}=\{\bm{x}\in\mathcal{C}\colon x_{j}>a\}\ .$
We solve the optimal spitting feature $j^{*}$ and splitting point $a^{*}$ from following optimization
@@
From Theorem 1, we can distinguish augmented class from known classes in node $\mathcal{C}$ by solving Eqn. (4), as it is convergent to optimizing the squared loss as training data increases. From a theoretical view, we require sufficient training data in each node $\mathcal{C}$, i.e., $n_{\mathcal{C},l}\geq\gamma n_{l}$ and $n_{\mathcal{C},u}\geq\gamma n_{u}$ for some $\gamma\in(0,1)$ in Theorem 1. Hence, we should construct relatively shallow tree models for theoretical guarantees.
On the other hand, we should take deeper tree models for better performance, and it is well-known that better predictions could be obtained by deepening tree models sufficiently Quinlan (1986); Hastie et al. (2009). From such recognition, our algorithm includes two main steps as follows:
Step-I: Exploration of augmented instances
@@
  • Solve the optimal splitting feature $j^{*}\in\mathcal{S}$ and point $a^{*}$ from the problem of Eqn. (4) under the constraints
@@
    $\min(n_{\mathcal{C}^{l}_{j,a},u},n_{\mathcal{C}^{r}_{j,a},u})$$\geq$ $\gamma n_{u}\ ,$ (6)
    which ensures sufficient instances in each child node for theoretical considerations;
@@
We stop the above procedure if there is no feasible solution for Eqn. (4) under constraints in Eqns. (5) and (6).
@@
where $\vartheta_{\mathcal{C}^{\prime}_{j},\kappa+1}(S_{l},S_{u})$ is an unbiased estimator for the proportion of augmented instances in $\mathcal{C}^{\prime}_{j}$ as shown by lemma 1.
Step-II: Improvement of prediction performance
@@
Given $m$ trees $\mathcal{T}^{\prime}_{1},\cdots,\mathcal{T}^{\prime}_{m}$, we can calculate the average augmented score of instance $\bm{x}\in\mathcal{X}$ by
$\mathcal{LF}^{\prime}_{\kappa+1}(\bm{x})=\frac{1}{m}\sum_{i\in[m]}\mathcal{T}^{\prime}_{i,\kappa+1}(\bm{x})\ ,$ (7)
and then select $\lfloor\theta n_{u}\rfloor$ instances in $S_{u}$ of the highest average augmented scores, and label them with pseudo-labels of augmented class. Without loss of generality, we denote by
@@
We stop the above procedure if all instances have the same label in a leaf, and get the final random tree $\mathcal{T}_{1}$. Similarly, we repeat the above procedure for $\mathcal{T}^{\prime}_{2},\cdots,\mathcal{T}^{\prime}_{m}$.
Getting $m$ random trees $\mathcal{T}_{1},\cdots,\mathcal{T}_{m}$, we predict the final label of instance $\bm{x}\in\mathcal{X}$ as
@@
where $\mathcal{T}_{i,k}(\bm{x})$ is the probability of the $k$-th category of tree $\mathcal{T}_{i}$ with leaves $\mathcal{C}_{i,1},\mathcal{C}_{i,2},\cdots,\mathcal{C}_{i,t_{i}}$, and is given by
$\sum_{j\in[t_{i}]}\frac{\mathbb{I}[\bm{x}\in\mathcal{C}_{i,j}]\sum_{(\bm{x}^{\prime},y)\in S_{l}\cup\tilde{S}}\mathbb{I}[\bm{x}^{\prime}\in\mathcal{C}_{i,j},y=k]}{\max(1,\sum_{(\bm{x}^{\prime},y)\in S_{l}\cup\tilde{S}}\mathbb{I}[\bm{x}^{\prime}\in\mathcal{C}_{i,j}])}\ .$
@@ 4 Deep Neural LACForest
Based on our augmented Gini impurity, we further explore deep neural forests for learning with augmented class, which presents powerful representations with an end-to-end training manner for complex data such as images and texts. Another advantage for neural forests is to take soft splits rather than hard splits of decision trees, which could yield smoother decision boundaries and avoid overfitting Kontschieder et al. (2015). Previous approaches generally focus on all known classes Kontschieder et al. (2015); Tanno et al. (2019); Ji et al. (2020); Li et al. (2024), while this work takes one step on learning with augmented class via deep neural forests.

### v2

@@ Abstract
Decision trees and forests have achieved successes in various real applications, most working with all testing classes known in training data. In this work, we focus on learning with augmented class via forests, where an augmented class may appear in testing data yet not in training data. We incorporate information of augmented class into trees’ splitting, that is, augmented Gini impurity, a new splitting criterion is introduced to exploit some unlabeled data from testing distribution. We then develop the Learning with Augmented Class via Forests (short for LACForest) approach, which constructs shallow forests according to the augmented Gini impurity and then splits forests with pseudo-labeled augmented instances for better performance. We also develop deep neural forests via an optimization objective based on our augmented Gini impurity, which essentially utilizes the representation power of neural networks for forests. Theoretically, we present the convergence analysis for our augmented Gini impurity, and we finally conduct experiments to evaluate our approaches. The code is available at https://github.com/nju-xuf/LACForest.
@@ 1 Introduction
How to handle distribution changes has become an important problem in a non-stationary learning environment Zhou (2022), and recent years have witnessed increasing attentions with various applications Gama et al. (2014); Geng et al. (2021); Wang et al. (2024). This work focuses on learning with augmented class, that is, the class distribution changes and an augmented class unseen in the training data may emerge during the testing process Da et al. (2014). Here, we take object recognition in autonomous driving for example: new objects may appear on roads yet not in the historical labeled data, and a reliable learning system should make good predictions over both known classes and augmented class.
Various approaches have been developed for learning with augmented class. Da et al. Da et al. (2014) studied decision boundaries for augmented class under the low-density assumption. Bendale and Boult Bendale and Boult (2016) and Rudd et al. Rudd et al. (2017) estimated the probability of an instance belonging to augmented class on the basis of extreme value theory. Mendes Júnior et al. Mendes Júnior et al. (2017) detected augmented class by a nearest neighbor method, and Liu et al. Liu et al. (2018) presented PAC guarantees for the detection of augmented class. Some generative networks have also been applied to learn augmented class Ge et al. (2017); Neal et al. (2018); Chen et al. (2021). Zhang et al. Zhang et al. (2020) gave an unbiased risk estimation by exploiting unlabeled data, and Shu et al. Shu et al. (2023) generalized such approach to arbitrary loss functions.
Decision trees and forests have achieved great successes in various applications with strong generalization in handling discrete features and exploring local regions Cutler et al. (2007); Qi (2012); Grinsztajn et al. (2022); Costa and Pedreira (2023). Most previous studies worked on the same distribution between training and testing data. For augmented class, Mu et al. Mu et al. (2017) and Liu et al. Liu et al. (2018) only studied its detection from labeled data of known classes, whereas it remains open on how to construct trees for learning with augmented class by exploiting unlabeled data from testing data.
This work aims to incorporate some useful information of augmented class into the construction of forests, and our main contributions can be summarized as follows:
  • We introduce a new splitting criterion, i.e., augmented Gini impurity, to incorporate information of augmented class from unlabeled data during the trees’ splitting. We develop the LACForest approach for learning with augmented class, which constructs shallow forests by augmented Gini impurity and considers pseudo-labeled augmented instances for better performance.
@@
  • From a theoretical view, we present the convergence analysis of our augmented Gini impurity with respect to both decision trees and deep neural trees. We finally conduct extensive experiments to validate the effectiveness of our proposed approaches and perform some parameter influence analysis.
The rest of this work is organized as follows: Section 2 gives some preliminaries. Section 3 introduces augmented Gini impurity and LACForest approach. Section 4 presents our deep neural approach. Section 5 conducts some extensive experiments. Section 6 concludes with future works.
@@ 2 Preliminaries
We focus on the semi-supervised setting for learning with augmented class Da et al. (2014); Liu et al. (2018); Zhang et al. (2020); Shu et al. (2023), and the goal is to learn a function $f\colon\mathcal{X}\rightarrow\mathcal{Y}$ from labeled data $S_{l}$ and unlabeled data $S_{u}$ with
$S_{l}=\{(\bm{x}^{l}_{1},y_{1}),\ldots,(\bm{x}^{l}_{n_{l}},y_{n_{l}})\}\text{ and }S_{u}=\{\bm{x}^{u}_{1},\ldots,\bm{x}^{u}_{n_{u}}\}\ .$
Here, each labeled example $(\bm{x}^{l}_{i},y_{i})$ is drawn i.i.d. from $\mathcal{D}_{\text{kc}}$, and each unlabeled instance $\bm{x}^{u}_{j}$ is sampled i.i.d. from $\mathcal{D}_{\mathcal{X}}$.
Let $\mathbb{I}[\cdot]$ be the indicator function, which returns $1$ if the argument is true and 00 otherwise. For integer $k>0$, we write $[k]=\{1,2,\cdots,k\}$, and denote by $\Delta_{k}$ the $k$-dimensional simplex. For $z\in\mathbb{R}$, we denote by $\lfloor z\rfloor$ the largest integer no more than $z$, and define $(z)_{+}=\max(0,z)$. Let $|A|$ be the cardinality of set $A$.
@@ 3 Our LACForest Approach
This section proposes the LACForest approach for learning with augmented class based on random forests, and the core idea is to introduce a new splitting criterion, augmented Gini impurity, to incorporate potential information of augmented instances during the construction of decision trees.
@@ 3.1 Augmented Gini Impurity
Denote by $n_{\mathcal{C},l}=|S_{\mathcal{C},l}|$ and $n_{\mathcal{C},u}=|S_{\mathcal{C},u}|$. We could define the augmented Gini impurity as a splitting criterion.
@@
$\vartheta_{\mathcal{C},k}(S_{l},S_{u})=\sum_{(\bm{x},y)\in S_{l}}(1-\vartheta_{\mathcal{C},\kappa+1}(S_{l},S_{u}))\frac{\mathbb{I}[\bm{x}\in\mathcal{C},y=k]}{\max(1,n_{\mathcal{C},l})}\ .$
Figure 1: An intuitive illustration on the difference between our criterion and original Gini impurity over a 2-dimensional dataset, by considering augmented class.
This definition essentially follows the original Gini impurity Breiman (1984), except with an additional term $\vartheta_{\mathcal{C},\kappa+1}(S_{l},S_{u})$, which aims to incorporate some information of augmented class into impurity measure.
Figure 1 presents an intuitive illustration on the difference between our augmented and original Gini impurity over a 2-dimensional synthetic dataset. It is observable that our augmented Gini impurity could properly take augmented data into consideration compared with Gini impurity.
For $\vartheta_{\mathcal{C},\kappa+1}(S_{l},S_{u})$, we also have
@@
For $\delta\in(0,1)$ and instance space $\mathcal{C}\subseteq\mathcal{X}$, the following holds with probability at least $1-\delta$ over $S_{l}$ and $S_{u}$
@@
where $\tilde{\bm{y}}\in\mathbb{R}^{\kappa+1}$ is the one-hot encoding of $y\in[\kappa+1]$. Intuitively, $\mathcal{L}^{*}_{\mathcal{C}}$ shows the minimum expected squared loss under the same prediction for all instances in space $\mathcal{C}$.
Based on Lemma 1 and Assumption 1, we have
@@
For $\delta\in(0,1)$ and instance space $\mathcal{C}\subseteq\mathcal{X}$, the following holds with probability at least $1-\delta$ over $S_{l}$ and $S_{u}$
@@
This theorem shows that our augmented Gini impurity can be viewed as a good estimation of $\mathcal{L}^{*}_{\mathcal{C}}$ with the convergence rate $O(1/\sqrt{\min\{n_{l},n_{u}\}})$. The detailed proof is given in Appendix A.2, which involves the equivalence between splitting criterions and loss functions, as well as the convergence of $\vartheta_{\mathcal{C},\kappa+1}(S_{l},S_{u})$ in Lemma 1.
@@ 3.2 Construction of LACForest
A tree model can be constructed by partitioning instance space into disjoint rectangular nodes recursively. Given node $\mathcal{C}\subseteq\mathcal{X}$, splitting feature $j\in[d]$ and splitting point $a\in\mathbb{R}$, we denote by the left and right children as
$\mathcal{C}^{l}_{j,a}=\{\bm{x}\in\mathcal{C}\colon x_{j}\leq a\}\ \text{ and }\ \mathcal{C}^{r}_{j,a}=\{\bm{x}\in\mathcal{C}\colon x_{j}>a\}\ ,$
respectively. We then solve the optimal splitting feature $j^{*}$ and splitting point $a^{*}$ from the following optimization
@@
From Theorem 1, we can distinguish augmented class from known classes in node $\mathcal{C}$ by solving Eqn. (4), as it converges to optimizing the squared loss as training data increases. From a theoretical view, we require sufficient training data in each node $\mathcal{C}$, i.e., $n_{\mathcal{C},l}\geq\gamma n_{l}$ and $n_{\mathcal{C},u}\geq\gamma n_{u}$ for some $\gamma\in(0,1)$ in Theorem 1; therefore, we should construct relatively shallow tree models.
On the other hand, we should take deeper tree models for better performance, and it is well-known that better predictions could be obtained by deepening tree models sufficiently Quinlan (1986); Hastie et al. (2009). Hence, our algorithm includes two main steps as follows:
3.2.1 Step-I: Exploration of Augmented Instances
@@
  • Solve the optimal splitting feature $j^{*}\in\mathcal{S}$ and point $a^{*}$ according to Eqn. (4) under the constraints
@@
    $\min(n_{\mathcal{C}^{l}_{j,a},u},n_{\mathcal{C}^{r}_{j,a},u})$$\geq$ $\gamma n_{u}\ .$ (6)
    This ensures sufficient instances in each splitting node from Theorem 1;
@@
The above procedure is stopped if there is no feasible solution for Eqn. (4) under constraints in Eqns. (5) and (6).
@@
where $\vartheta_{\mathcal{C}^{\prime}_{j},\kappa+1}(S_{l},S_{u})$ is an unbiased estimator for the proportion of augmented instances in $\mathcal{C}^{\prime}_{j}$, given by Definition 1.
3.2.2 Step-II: Improvement of Prediction Performance
@@
Given $m$ trees $\mathcal{T}^{\prime}_{1},\cdots,\mathcal{T}^{\prime}_{m}$, we calculate the average augmented score of instance $\bm{x}\in\mathcal{X}$ by
$\mathcal{LF}^{\prime}_{\kappa+1}(\bm{x})=\frac{1}{m}\sum_{i\in[m]}\mathcal{T}^{\prime}_{i,\kappa+1}(\bm{x})\ .$ (7)
We then select $\lfloor\theta n_{u}\rfloor$ instances in $S_{u}$ of the highest average augmented scores, and label them with pseudo-labels of augmented class. Without loss of generality, we denote by
@@
The above procedure is stopped if all instances have the same label in a leaf, and get the final random tree $\mathcal{T}_{1}$. Similarly, we repeat the above procedure for $\mathcal{T}^{\prime}_{2},\cdots,\mathcal{T}^{\prime}_{m}$.
Given $m$ random trees $\mathcal{T}_{1},\cdots,\mathcal{T}_{m}$, we predict the final label of instance $\bm{x}\in\mathcal{X}$ as
@@
where $\mathcal{T}_{i,k}(\bm{x})$ is the probability of the $k$-th category of tree $\mathcal{T}_{i}$ with leaves $\mathcal{C}_{i,1},\mathcal{C}_{i,2},\cdots,\mathcal{C}_{i,t_{i}}$, i.e.,
$\mathcal{T}_{i,k}(\bm{x})=\sum_{j\in[t_{i}]}\frac{\mathbb{I}[\bm{x}\in\mathcal{C}_{i,j}]\sum_{(\bm{x}^{\prime},y)\in S_{l}\cup\tilde{S}}\mathbb{I}[\bm{x}^{\prime}\in\mathcal{C}_{i,j},y=k]}{\max(1,\sum_{(\bm{x}^{\prime},y)\in S_{l}\cup\tilde{S}}\mathbb{I}[\bm{x}^{\prime}\in\mathcal{C}_{i,j}])}\ .$
@@ 4 Deep Neural LACForest
This section explores deep neural forests via augmented Gini impurity for learning with augmented class, which presents powerful representations with an end-to-end training manner for complex data such as images and texts. For neural forests, another advantage is to take soft splits rather than hard splits of decision trees, which could yield smoother decision boundaries and avoid overfitting Kontschieder et al. (2015). In this work ,we take one step on learning with augmented class via deep neural forests, while previous works mostly focused on all known classes Kontschieder et al. (2015); Tanno et al. (2019); Ji et al. (2020); Li et al. (2024).